# Flows of vector fields and the Kalman Theorem

### v1

@@ Abstract.
We investigate the relations between the Kalman Theorem and the Chow-Rashevskiĭ Theorem or, more precisely, the general theory of flows tangent to non-integrable distributions. The main results consist of two proofs of the Kalman Theorem, which are alternative to the most common ones and which enlighten the above relations and stimulate generalisations in various directions.
@@ 1. Introduction
for some non zero constant matrices $A=(A^{i}_{j})$ and $B=(B^{i}_{a})$ and in which the controls $u(t)$ are measurable functions taking values in a bounded set $\mathscr{K}\subset\mathbb{R}^{m}$. For a fixed $\overline{q}\in\mathscr{Q}$, denote by $\mathcal{C}_{\overline{q}}$ the set of all points $q\in\mathscr{Q}$ that are controllable to $\overline{q}$, i.e. for which there is a control $u(t)$ which determines a solution $q(t)$ to (1.1) that starts from $q$ and reaches $\overline{q}$ in a finite time. The famous Kalman Theorem is an important criterion to determine whether the controllable points to the origin $0_{\mathbb{R}^{n}}$ contains an open set. It states that
@@
The proof can be found in many textbooks (as for instance [1, 3]), and it essentially uses just results in Linear Algebra and Convex Geometry.
This maximal dimension condition is very similar to the condition on the Lie bracket generated spaces, which appear in the corollary of the Chow-Rashevskiĭ Theorem concerning the local controllability of systems of the form $\dot{q}=\sum_{a=0}^{m}X_{a}(q)u^{a}$. In fact, the relation between those theorems is much more than a mere likeness, since
  • (a)
    The vector space $V$ generated by the vectors (1.2) coincides with the space that is generated by the Lie brackets at $q=0$ of an appropriate collection of $m+1$ vector fields $X_{a}:\mathscr{Q}\to\mathbb{R}^{n}$ (and this is exactly the space considered in the statement of the Chow-Rashevskiĭ Theorem (see §3.2 below)), namely of a collection of vector fields, among which the components of first vector field $X_{0}$ are linear functions of the coordinates of $\mathscr{Q}$, while the components of all the others are constant functions;
  • (b)
    The sufficiency parts of both results, i.e. of the Kalman Theorem and of the mentioned corollary of the Chow-Rashevskiĭ  Theorem, can be derived from a very general criterion by Sussmann ([13]; see also [3, Thm. 3.29 & Ex.3.3]);
  • (c)
    A couple of results by Hermann [6] and Nagano [10] provide a common argument for proving the necessity parts of both theorems (also on this regard, see e.g. [1, 3]).
Motivated by these facts, there is a common belief that the Kalman Theorem is essentially equivalent to the above quoted corollary of the Chow-Rashevskiĭ Theorem. Since it is not hard to check that “the Kalman Theorem yields such a corollary of the Chow-Rashevskiĭ Theorem” (see e.g. §4 below), the existence or not of an equivalence between those two theorems depends on whether or not the Chow-Rashevskiĭ Theorem implies the Kalman Theorem in a direct way. The aim of this paper is precisely to clarify to which extent such an implication exists and hence whether the above expected equivalence is actually correct.
@@
Our achievements consist of two (at best of our knowledge, new) proofs for the Kalman Theorem that answer the above question and that we believe can be used also to determine convenient generalisations of the Kalman Theorem. The first proof is short, direct and makes fully explicit a detailed relation between the (above described corollary of the) Chow-Rashevskiĭ  Theorem and the Kalman Theorem. It shows that the first implies the second on the basis of the following simple (but crucial) fact: if $\mathscr{K}$ contains a neighbourhood of the origin – and thus, in particular, a convex symmetric neighbourhood of the $0_{\mathbb{R}^{n}}$ – then also the controllable set $\mathcal{C}_{q=0_{\mathbb{R}^{n}}}$ of (1.1) contains a convex set (see (1) of Proposition 5.1, below). We guess that this new proof admits generalisation for other kinds of (possibly non-linear) control systems, provided that similar convexity properties for the controllable sets can be established. On this regard, we would like to mention that investigations, involving ideas very close to those of our first proof, can be found in [8], where interesting extensions of Kalman criterion are reached.
@@
Our second proof is much longer than the first, it is based on a discussion on the higher dimensional space =×$\mathscr{M}=\mathbb{R}\times\mathscr{Q}\times\mathscr{K}$ and it does not involve in a direct way the Chow-Rashevskiĭ  Theorem – it just uses a few results on flows, which are properties that can be taken as ancestors of the Chow-Rashevskiĭ  Theorem. However, these drawbacks are balanced by the fact that our second proof makes no use of any information on convexity properties of the controllable sets. This makes this second proof even more appropriate than the first for generalisations to non-linear control systems. In fact, generalisations of this kind have been established in [5], where new local controllability criteria are given and diverse controllability problems, for which all previously known criteria are inconclusive, are solved.
@@
The paper is structured as follows. After section §2, in which we give a few preliminaries and introduce some convenient notation, in §3 we recall the statements of the Chow-Rashevskiĭ and the Kalman Theorem, on which our discussion is based. In §4 we prove that, for the collections of $m+1$ vector fields $X_{\alpha}$ of the above described form, the property determined by the Chow-Rashevskiĭ Theorem can be inferred directly from the Kalman Theorem. In §5 and §6 there are our main results, i.e. the mentioned couple of new proofs for the Kalman Theorem, one based on the Chow-Rashevskiĭ Theorem and the other on general properties of flows.
@@ 2.2. Pushed-forward vector fields and their flows
The following fact is well known (see e.g. [14, Prop. 1.55]): For any two $\mathcal{C}^{k}$ vector fields $X$, $Y$, $k\geq 1$, on $\mathscr{U}$ and any $\mathcal{C}^{2}$ diffeomorphism $\varphi:\mathscr{U}\to\mathscr{V}\subset\mathbb{R}^{n}$ then $\varphi_{*}([X,Y])=\left[\varphi_{*}(X),\varphi_{*}(Y)\right]$. This and a direct calculation imply that the following relation between flows holds at all points where both sides are defined:
@@ 3.1. Orbits of sets of vector fields and the Chow-Rashevskiĭ Theorem
This notation being fixed, we can state Sussmann’s refined and improved version of the Chow-Rashevskiĭ Theorem as follows (see also [4, 5]).
Theorem 3.1 (Chow-Rashevskiĭ-Sussmann [2, 11, 12]).
@@ 5. A proof of the Kalman Theorem, based on the Chow-Rashewskiĭ Theorem
(1) is a well known property of linear system, which can be easily checked by just considering the classical integral representation formula for the solutions to (1.1) (see e.g. [9, §II.2]).

### v2

@@ Abstract.
We give two proofs of the Kalman Theorem, alternative to the most common ones, which infer such a classical result of Control Theory using just very basic facts on flows of vector fields. These proofs are apt to be generalised in diverse directions – in fact one of them has been already generalised, yielding new criteria for local controllability of non-linear real analytic controlled systems.
@@ 1. Introduction
for some non zero constant matrices $A=(A^{i}_{j})$ and $B=(B^{i}_{a})$ and in which the controls $u(t)$ are measurable functions taking values in a bounded set $\mathscr{K}\subset\mathbb{R}^{m}$. For a fixed $\overline{q}\in\mathscr{Q}$, denote by $\mathcal{C}_{\overline{q}}$ the set of all points $q\in\mathscr{Q}$ that are controllable to $\overline{q}$, i.e. for which there is a control $u(t)$ which determines a solution $q(t)$ to (1.1) that starts from $q$ and reaches $\overline{q}$ in a finite time. A local version of the famous Kalman Theorem consists of a criterion for determining whether the set $\mathcal{C}_{\overline{q}=0_{\mathbb{R}^{n}}}$ of controllable points to the origin contains an open set. It can be stated as follows.
@@
The proof can be found in many textbooks (as for instance [1, 3]), and it essentially uses results in Linear Algebra and Convex Geometry.
The maximal dimension condition of the Kalman Theorem is very much reminiscent of a known corollary of the Chow-Rashevskiĭ Theorem on the local controllability of systems of the form $\dot{q}=\sum_{a=0}^{m}X_{a}(q)u^{a}$. In fact, the sufficiency parts of the Kalman Theorem and of the mentioned corollary of the Chow-Rashevskiĭ  Theorem, can be derived from a very general criterion by Sussmann ([18] or [3, Thm. 3.29 Ex.3.3]; on this regard, see also foundational Sussmann’s and Stefan’s papers [16, 17, 13, 14]). Furthermore, a couple of results by Hermann [7] and Nagano [11] provide a common argument for proving the necessity parts of both theorems (on this regard, see also [1, 3]). Due to all this, it is commonly understood that the Kalman Theorem is equivalent to the above quoted corollary of the Chow-Rashevskiĭ Theorem. Since it is not hard to check that “the Kalman Theorem yields such a corollary of the Chow-Rashevskiĭ Theorem” (see e.g. §4 below for a short proof or [4, 15] for proofs holding in much more general settings), on the basis of the above common belief it is sensible to expect the existence of a straightforward way of deriving the Kalman Theorem from the Chow-Rashevskiĭ Theorem. In this paper we make explicit such a direct implication in two ways. More precisely, our main achievements consist of two (at best of our knowledge, new) proofs of the Kalman Theorem, which also appear to be appropriate for generalisations to fully nonlinear control systems. Actually, the second of our proofs have been successfully exploited in [6] and have led to new controllability criteria for nonlinear real analytic control system.
@@
Our first proof is short and gives a simple way to pass from the (above described corollary of the) Chow-Rashevskiĭ  Theorem to the Kalman Theorem. It shows that the first implies the second just on the basis of the following simple (but crucial) fact: if $\mathscr{K}$ contains a neighbourhood of the origin – and thus, in particular, a convex symmetric neighbourhood of the $0_{\mathbb{R}^{n}}$ – then also the controllable set $\mathcal{C}_{q=0_{\mathbb{R}^{n}}}$ of (1.1) contains a convex set (see (1) of Proposition 5.1, below). As we mentioned above we expect that this new proof admits generalisations for other kinds of control systems, provided that similar convexity properties for the controllable sets can be established. On this regard, we would like to mention that investigations, involving ideas very close to those of our first proof, can be found in [9], where interesting extensions of the Kalman criterion are reached.
@@
Our second proof is much longer than the first, it is based on a discussion on the higher dimensional space =×$\mathscr{M}=\mathbb{R}\times\mathscr{Q}\times\mathscr{K}$ and it does not involve in a direct way the Chow-Rashevskiĭ  Theorem – it just uses a few results on flows, which are properties that can be taken as ancestors of the Chow-Rashevskiĭ  Theorem. However, these drawbacks are balanced by the fact that our second proof makes no use of any information on convexity properties of the controllable sets. This makes this second proof even more appropriate than the first for generalisations to non-linear control systems. Indeed, such generalisations have been established in [6], where new local controllability criteria are given and diverse controllability problems, for which all previously known criteria are inconclusive, are solved.
@@
The paper is structured as follows. After section §2, in which we give a few preliminaries and introduce some convenient notation, in §3 we recall the statements of the Chow-Rashevskiĭ and the Kalman Theorem, on which our discussion is based. In §4, for the sake of completeness, we exhibit a direct proof that, for the collections of $m+1$ vector fields $X_{\alpha}$ of the above described form, the property determined by the Chow-Rashevskiĭ Theorem can be inferred directly from the Kalman Theorem (as we mentioned above, other proofs that hold in a much more general setting are given in [4, 15]). In §5 and §6 there are our main results, i.e. the mentioned new proofs for the Kalman Theorem, one based on the Chow-Rashevskiĭ Theorem and the other on general properties of flows.
@@
Acknowledgments. We are sincerely grateful to David Chillingworth for kindly pointing out Peter Štefan’s works to us.
@@ 2.2. Pushed-forward vector fields and their flows
The following fact is well known (see e.g. [19, Prop. 1.55]): For any two $\mathcal{C}^{k}$ vector fields $X$, $Y$, $k\geq 1$, on $\mathscr{U}$ and any $\mathcal{C}^{2}$ diffeomorphism $\varphi:\mathscr{U}\to\mathscr{V}\subset\mathbb{R}^{n}$ then $\varphi_{*}([X,Y])=\left[\varphi_{*}(X),\varphi_{*}(Y)\right]$. This and a direct calculation imply that the following relation between flows holds at all points where both sides are defined:
@@ 3.1. Orbits of sets of vector fields and the Chow-Rashevskiĭ Theorem
This notation being fixed, we can state Sussmann’s refined and improved version of the Chow-Rashevskiĭ Theorem as follows (see also [13, 14] for an even more general variant of Sussmann’s version and [5, 6] for discussions and improvements in various directions).
Theorem 3.1 (Chow-Rashevskiĭ-Sussmann [2, 12, 16, 17]).
@@ 5. A proof of the Kalman Theorem, based on the Chow-Rashewskiĭ Theorem
(1) is a well known property of linear system, which can be easily checked by just considering the classical integral representation formula for the solutions to (1.1) (see e.g. [10, §II.2]).